OpenStudy (anonymous):

if f(a) = a-2 and F(a,b) = b^2 + a, then F(3, f(4)) is ?

OpenStudy (anonymous):

Ok so at first you should find f(4) using f(a) = a-2. post your result and I will help you through it

OpenStudy (anonymous):

so is f(4) = 2

OpenStudy (anonymous):

yes, Now using F(a,b) = b^2; if we want to find F(3,f(4)). what shall we do?

OpenStudy (anonymous):

isn't F(a,b) = b^2 + a

OpenStudy (anonymous):

my bad. that doesn't make a big difference though, use the value we found out in f(4) as b, square it and add it with a (which is 3). that's it

OpenStudy (anonymous):

thank so much. i finally know how to it.

OpenStudy (anonymous):

no problem